The travel-to-work time for residents of the 22 largest cities in the United States is reported in the
2003 Information Please Almanac. Suppose that a preliminary simple random sample of residents
of San Francisco is used to develop a planning value of 5.95 minutes for the population standard
deviation.

a. If we want to estimate the population mean travel-to-work for San Francisco residents with a
margin of error of 2 minutes, what sample size should be used? Assume 95% confidence.
b. If we want to estimate the population mean travel-to-work for San Francisco residents with a
margin of error of 30 seconds, what sample size should be used? Assume 99% confidence.

Answer by stanbon(75887)

The travel-to-work time for residents of the 22 largest cities in the United States is reported in the 2003 Information Please Almanac. Suppose that a preliminary simple random sample of residents of San Francisco is used to develop a planning value of 5.95 minutes for the population standard
deviation.
a. If we want to estimate the population mean travel-to-work for San Francisco residents with a margin of error of 2 minutes, what sample size should be used? Assume 95% confidence.
Since E = z*s/sqrt(n), n = [z*s/E]^2 = [1.96*5.95/2]^2 = 34
------------------------------
b. If we want to estimate the population mean travel-to-work for San Francisco residents with a margin of error of 30 seconds, what sample size should be used? Assume 99% confidence.
n = [2.5758*5.95/(1/2)]^2 = 940 when rounded up
-------------
Cheers,
Stan H.
----------------